### v1

# Explain with Visual Key-Points Like a Real Mentor ! A Benchmark for Visual Solution-Explaining

Abstract
With the rapid advancement of mathematical reasoning capabilities in large language models (LLMs), AI systems are increasingly being adopted in educational settings to support students’ comprehension of problem-solving processes. However, a critical component remains underexplored in current LLM-generated explanations: visual explanation. In real-world instructional contexts, human tutors routinely employ visual aids—such as diagrams, markings, and highlights—to enhance conceptual clarity. To bridge this gap, we introduce a novel task of visual solution explanation, which requires not only solving problems but also generating explanations that incorporate newly introduced visual elements essential for understanding (e.g., auxiliary lines, annotations, or geometric constructions). To evaluate model performance on this task, we propose MathExplain, a multimodal benchmark consisting of 997 math problems annotated with visual keypoints and corresponding explanatory text that references those elements. Our empirical results show that while some closed-source models demonstrate promising capabilities on visual solution-explaining, current open-source general-purpose models perform inconsistently, particularly in identifying relevant visual components and producing coherent keypoint-based explanations. We expect that visual solution-explaining and the MathExplain dataset will catalyze further research on multimodal LLMs in education and advance their deployment as effective, explanation-oriented AI tutors. Code and data will be released publicly.
1 Introduction
The traditional one-to-many educational model (i.e., one teacher for multiple students) is gradually transitioning to one-to-one personalized tutoring systems and online learning (Mukul & Büyüközkan, 2023). Recent developments in Multimodal Large Language Models (MLLMs) have opened new opportunities for effective learning, such as estimating question difficulty for recommendation systems (Park et al., 2024), assisting teachers in curriculum planning (Hu et al., 2024), and supporting interactive tutoring systems (Chevalier et al., 2024). In particular, numerous studies (Liu et al., 2023; Uesato et al., 2022; Lu et al., 2023) have focused on enhancing the mathematical reasoning abilities of MLLMs. As a result, MLLMs have led many students to use them as tools when faced with mathematical questions (Pardos & Bhandari, 2024).
However, from a student’s perspective, relying solely on the reasoning footprints of MLLMs may not always be the best way to understand problems (Pardos & Bhandari, 2024; Jia et al., 2024). One might wonder what distinguishes a broadly comprehensible explanation from a solution that merely yields the correct answer—for either a human or a model? A critical factor is the use of visual cues.
In actual educational settings, Dual Coding Theory (DCT) naturally occurs, providing effective learning opportunities for students (Paivio, 2013; 1990). According to DCT, combining verbal and visual information enhances student comprehension (Clark & Paivio, 1991). As illustrated in Figure 1, human mentors often use visual scaffolding—such as annotated diagrams or highlighted keypoints on a blackboard—to foster intuitive understanding (Arcavi, 2003; Stylianou, 2010; Lee et al., 2025). In contrast, current AI models lack the capacity to generate such visual explanations. Moreover, existing datasets for training and evaluation are insufficient for developing models capable of producing these forms of multimodal instructional support, as summarized in Table 1.
Figure 1: A student solving a math problem often benefits from visual cues—such as lines, symbols, or highlights—that human instructors use to aid understanding, unlike current AI models that focus solely on textual solutions. To serve as effective educational assistants, machines must go beyond answer generation and emulate human-like explanation strategies by explicitly incorporating and referencing visual elements.
Benchmark ${Text}_{problem}$ ${Image}_{problem}$ ${Text}_{solution}$ ${Image}_{solution}$ Task
MATH (Hendrycks et al., 2021) Problem-Solving
MathVista (Lu et al., 2023) Problem-Solving
MATH-Vision (Wang et al., 2024) Problem-Solving
MathExplain (Ours) Solution-Explaining
Table 1: MathExplain is designed to evaluate the capacity of models to generate visually grounded explanations of mathematical solutions, in contrast to existing benchmarks that primarily assess answer correctness.
To address these limitations, we introduce visual solution-explaining, a novel task designed to evaluate and advance models’ ability to generate educationally effective, visually grounded mathematical explanations. To benchmark performance on visual solution-explaining, we propose multimodal Mathematics Explain (MathExplain) benchmark. The visual solution-explaining task requires identifying visual keypoints—elements not present in the original problem but crucial for understanding (e.g., lines, angles, annotations)—and generating explanatory text that explicitly refers to them. The MathExplain includes not only the original problem images and text but also solution images, explanatory text, and annotations of the visual keypoints that illustrate how the solution images differ from the originals (see Table 1). By providing textual representations of these visual keypoints, the benchmark allows even models—which currently struggle to generate fine-grained figures (Wu et al., 2024; Jo et al., 2024)—to capture and convey visually salient details in textual form. Notably, we emphasize the educational value of MathExplain, as it is derived from real-world instructional contexts.
Experiments on MathExplain demonstrate that, while closed-source models show potential on visual solution-explaining, current open-source general models still struggle to reliably identify visual keypoints. Even mathematical models, which are fine-tuned on mathematical domains, remain primarily focused on problem-solving and often do not produce sufficiently structured outputs to highlight critical visual details. Nonetheless, we believe MathExplain will spark further research into effectively harnessing AI models for educational applications and foster progress in problem-solving capabilities.
Our contributions are as follows:
  1. 1.
    A visual solution-explaining task that supports students’ educational comprehension by identifying critical visual keypoints and generating explanatory text that explicitly references them.
  2. 2.
    A MathExplain benchmark, rooted in authentic educational contexts, to rigorously assess visual solution-explaining performance and facilitate further research on model-based explanations in real-world settings.
  3. 3.
    Extensive experimental evaluations of state-of-the-art MLLMs on our visual solution-explaining task, highlighting current limitations in recognizing and leveraging crucial visual keypoints for effective personalized learning support.
2 Related Works
Large Language Models for Education.
Recent advances in Large Language Models (LLMs) (Brown et al., 2020) have sparked significant interest in educational applications, particularly in personalized learning and automated tutoring (Park et al., 2024; Chevalier et al., 2024). These systems can adapt to individual learners by providing tailored feedback and customized curriculum (Hu et al., 2024; Macina et al., 2023; Feng et al., 2023). Beyond textual analysis, educational research indicates that providing both textual and visual information can more effectively bolster comprehension and memory than text alone (Arcavi, 2003; Stylianou, 2010; Lee et al., 2025). As Clark & Paivio (1991) notes, this advantage stems from the use of dual representations, which offer multiple retrieval cues and foster richer mental models for learners. To further harness these multimodal elements and provide more comprehensive learning opportunities, we introduce the visual solution-explaining task for an MLLM equipped with advanced multimodal understanding capabilities. This task enables the model to autonomously identify crucial visual keypoints—essential for understanding visual information—and integrate them into its explanations. By doing so, model can deliver more comprehensive educational support, ultimately enhancing the overall quality of student learning experiences.
Multimodal Mathematical Benchmarks.
Current MLLMs excel at solving mathematical problems, making them highly useful tools for students (Zhuang et al., 2024; Luo et al., 2025). To evaluate these models, mathematical benchmarks (Cobbe et al., 2021; Hendrycks et al., 2021) have played a key role in assessing the reasoning capabilities of MLLMs. With the rise of multimodal models, benchmarks such as MathVista (Lu et al., 2023), Math-Vision (Wang et al., 2024), and MathVerse (Zhang et al., 2024) have extended these evaluations by introducing image-based math problems. Additionally, benchmarks like OlympiadBench (He et al., 2024) and MM-MATH (Sun et al., 2024) evaluate not only final answers but also the detailed solution process, providing deeper insights into model reasoning strategies. However, most existing benchmarks do not prioritize educational objectives. They generally focus on problem-solving performance rather than on the explanatory or pedagogical quality of solutions. Furthermore, the solutions they produce are predominantly textual, lacking the visual aids that can greatly enhance understanding. To address these gaps, we propose MathExplain, a new benchmark designed to assess a model’s visual solution-explaining capability—its ability to provide rich, multimodal explanations. By providing diverse and informative visuals, our benchmark advances educational research and supports the development of better teaching models.
3 MathExplain Benchmark
Figure 2: An overview of MathExplain. The MathExplain consists of multimodal problem–solution pairs and visual keypoints, generated through a Human–AI annotation.
Figure 3: We propose two subtasks to robustly analyze visual solution-explaining capacity: (1) Visual Keypoint Identification, which challenges machines to generate visual keypoints useful for future explanation, and (2) Keypoint-based Explanation Generation, which requires models to generate explanations that explicitly reference the identified visual keypoints.
MathExplain is a multimodal visual solution-explaining benchmark consisting of 997 instances. Each instances contains a problem text ($T_{p}$), a problem image ($I_{p}$), a explanatory solution text ($T_{s}$), a solution image ($I_{s}$), and visual keypoints ($VK$), which detail how the solution image differs from the original problem image. To create a benchmark that can assess the visual solution-explaining capabilities of MLLMs while considering their limitations in generating fine-grained figures, we define the visual keypoint $VK$ in Section 3.1. An overview of MathExplain and its construction is illustrated in Figure 2.
3.1 Benchmark Construction
In-house Data Preparation.
We extract 997 instances of multimodal problem–solution pairs $\langle T_{p},\,I_{p},\,T_{s},\,I_{s}\rangle$ from an in-house mathematics education platform. All instances were authored by domain experts in mathematics to support effective student learning. The problems cover topics ranging from middle to high school level, primarily focusing on geometry and graph interpretation. All benchmark data were carefully curated to ensure compliance with copyright regulations.
To benchmark the models’ ability to recognize visual keypoints, we ensure that each solution image $I_{s}$ is derived from the corresponding problem image $I_{p}$ by adding only new elements—such as points, angles, lines, regions, and symbols—while preserving the original structure. For instance, in Figure 2, symbols and dashed lines are added to problem image $I_{p}$ to produce solution image $I_{s}$. This setup allows us to accurately evaluate whether a model can identify the critical visual keypoints and effectively incorporate them into its explanation.
For robust evaluation, we restrict the dataset to single-image math problems with either multiple-choice or short-answer formats. We focus on the domains of geometry and mathematical functions to ensure that visual context is essential for solving each problem. Each sample in MathExplain consists of two natural language texts (the problem text $T_{p}$ and the solution text $T_{s}$) and two RGB images (the problem image $I_{p}$ and the solution image $I_{s}$).
Visual Keypoints Annotation Process.
To generate the visual keypoints (VK) in textual form, we use GPT-4o (Achiam et al., 2023) as an annotation tool. Identifying and formatting visual keypoints is a labor-intensive process; this tool assists human annotators by generating initial data frames. In our annotation workflow, the initial visual keypoints (VK) are generated by the AI tool and subsequently verified and refined by human annotators.
We define visual keypoint formatting guidelines to keep every visual keypoint at the same level, enabling precise evaluation of the model’s visual solution-explaining performance:
  1. (1)
    If any elements (points, lines, angles, regions, or symbols) are added or modified, they must be written in a {key: value} format. In this format, the key specifies the newly added or changed element, and the value describes how it is introduced by referencing surrounding elements.
  2. (2)
    Symbols serve to represent anything that cannot be expressed as points, lines, angles, or regions. Examples include parallel marks, equal-length marks, right-angle marks, and length-label marks.
Following these principles, a set of the visual keypoints $\{vk^{AI}_{1},\dots,vk^{AI}_{n}\}\in VK^{AI}$ are generated by AI tools. After the model-generated results are produced, four human annotators with undergraduate or master’s degrees in science or engineering assess the quality of $VK^{AI}$ and check whether it adheres to the formatting guidelines. They then refine the $VK^{AI}$ to ensure both high quality and consistency with these principles. Lastly, the AI annotator translates the benchmark from Korean to English, and two human annotators verify and refine the translated content, ultimately forming the $VK$. The final structure of MathExplain is represented as $\langle T_{p},I_{p},T_{s},I_{s},VK\rangle$.
Total problem–solution pairs 997
- Geometry 765
   - Multiple-choice 466
   - Short-answer questions 299
- Geometry 232
   - Multiple-choice questions 141
   - Short-answer questions 91.0
Average number of $VK$ 3.8
Maximum words in $T_{p}$ 211
Maximum words in $T_{s}$ 362
Maximum words in $k(I_{s})_{n}$ 45
Average words in $T_{p}$ 53.2
Average words in $T_{s}$ 102.4
Average words in $vk_{n}$ 12.1
Table 2: Statistics of the MathExplain benchmark size and length.
Figure 4: Visualization of subject-chapter diversity.
3.2 Data Analysis
The MathExplain benchmark consists of 997 problem–solution pairs: 765 (76.7%) geometry problems and 235 (23.5%) graph problems. Among these, 587 (58.8%) are multiple-choice questions and 423 (42.4%) are short-answer questions. It spans 17 chapters (see Figure 4) and 33 sections (see Appendix A). On average, each sample contains about 3.8 visual keypoints $VK$, derived from annotations. These keypoints fall into five main categories: points, lines, regions, and symbols. The symbol category is further divided into parallel marks, equal-length marks, right-angle marks, and length-label marks. Additional statistical details about $VK$, and length statistics for the problem text $T_{p}$, the solution text $T_{s}$, and the visual keypoint components $vk_{n}$ are provided in Table 2.
4 Task Definition
We propose two tasks to evaluate a model’s visual solution-explaining capability. visual solution-explaining requires identifying visual key-points useful for future explanation and generating explanations that explicitly reference them. These tasks offer a chance to assess how well the model understands multimodal educational elements and produces informative explanations along with visual keypoints. The two tasks are illustrated in Figure 3.
Visual Keypoint Identification.
The first task evaluates the model’s ability to infer visual keypoints that are crucial for comprehension. Given a problem image $I_{p}$ and problem text $T_{p}$, the model must infer the appropriate the visual keypoints $VK$ necessary for effective understanding. The $VK$ is represented as a list of dictionaries, where each key corresponds to a newly introduced visual element and the value provides a description of that element based on surrounding contextual features.
Keypoint-based Explanation Generation.
The second task evaluates whether the model can effectively generate explanatory text by appropriately using the visual keypoints it has already recognized. Given a problem consisting of an image $I_{p}$ and text $T_{p}$, along with the visual keypoints $VK$, the model is required to produce a solution explanation $T_{s}$ that references the relevant visual elements.
5 Experiments
Models
We evaluate both generalist and specialist MLLMs for multimodal mathematical problem solving. The generalist models include widely-used open-source MLLMs—LLaVA-1.6 (Liu et al., 2024), Molmo (Deitke et al., 2024), and Qwen-2.5-VL (Bai et al., 2025)—as well as proprietary models such as GPT-4o (Achiam et al., 2023) and Gemini 2.0 Flash (Google DeepMind, 2024). Specialist models focused on mathematics include Math-PUMA (Zhuang et al., 2024) and URSA (Luo et al., 2025). Details of the experimental setup are provided in Appendix A.
5.1 Visual Keypoint Identification
Given a problem image $I_{p}$ and accompanying text $T_{p}$, the model is expected to generate a set of visual keypoints $\{vk^{\text{Model}}_{1},\dots,vk^{\text{Model}}_{j}\}\in VK^{\text{Model}}$ that are essential for explaining the problem’s solution.
Metric
We evaluate the model’s output by comparing the generated set $VK^{\text{Model}}$ with the ground-truth set $VK^{\text{GT}}=\{vk^{\text{GT}}_{1},\dots,vk^{\text{GT}}_{i}\}$ using exact match (EM) at the individual keypoint level. We report set-level precision ($EM_{P}$), recall ($EM_{R}$), and F1-score ($EM_{F1}$) as follows:
$EM_{P}=\frac{|VK^{\text{Match}}|}{|VK^{\text{Model}}|},\quad EM_{R}=\frac{|VK^{\text{Match}}|}{|VK^{\text{GT}}|},\quad EM_{F1}=2\times\frac{EM_{P}\times EM_{R}}{EM_{P}+EM_{R}}$
To establish matching between the two sets, two steps are required: (1) parsing the natural language model output into a structured set of discrete keypoint descriptions, and (2) determining whether each predicted keypoint semantically matches a ground-truth point. For parsing, we employ GPT‑4o (Achiam et al., 2023) to convert the generated text into JSON format. Each keypoint $vk_{n}$ is represented with an "element" field indicating the added visual component and a "description" field explaining how it modifies or extends the original content. Semantic matching is also performed using GPT‑4o, leveraging its strong language understanding capabilities. This allows for flexible alignment between corresponding elements despite label variations (e.g., a ground-truth “dotted line AB” matching a model output labeled “AZ”), enabling robust evaluation while accounting for surface-level discrepancies.
Figure 5: Failure cases of Qwen-2.5-VL and Math-PUMA in the Visual Keypoint Identification task. Qwen-2.5-VL shows ability to extract keypoints, whereas Math-PUMA mainly focuses on problem-solving and fails to identify them.
Overall Geometry Graph
Model #Params $EM_{P}$ $EM_{R}$ $EM_{F1}$ $EM_{P}$ $EM_{R}$ $EM_{F1}$ $EM_{P}$ $EM_{R}$ $EM_{F1}$
OG Molmo 7B 0.188 0.111 0.121 0.185 0.107 0.117 0.197 0.125 0.133
LLaVA-1.6 7B 0.189 0.112 0.122 0.178 0.099 0.108 0.226 0.155 0.166
Qwen2.5-VL 7B 0.146 0.155 0.123 0.145 0.149 0.120 0.149 0.175 0.134
OM Math-PUMA 7B 0.038 0.058 0.036 0.029 0.058 0.032 0.067 0.057 0.050
URSA 8B 0.110 0.096 0.090 0.116 0.100 0.094 0.089 0.085 0.079
CG Gemini 2.0 Flash - 0.473 0.232 0.279 0.480 0.220 0.273 0.448 0.271 0.301
[…]
Table 3: Experimental results for the Visual Keypoint Identification task on MathExplain. The best scores are boldfaced, and the second-best scores are underlined.
Result
As shown in Table 3, none of the current Open-source General (OG) or Open-source Mathematical (OM) models effectively identify visual keypoints. In contrast, Closed-source General (CG) models—Gemini 2.0 Flash and GPT-4o—demonstrate significantly stronger performance across all metrics, although they still exhibit certain limitations. Notably, both CG and OG models have higher $EM_{P}$ than $EM_{R}$, implying that while they capture some intuitive keypoints, they struggle with more detailed or complex ones. Despite being designed for mathematical reasoning, OM models perform unexpectedly poorly, because they often fail to follow instructions properly (see Figure 5). These models, with their strong focus on problem-solving, are not suitable for generating insightful explanations for students. In contrast, OG models at least make an effort to follow instructions when identifying visual keypoints. Furthermore, CG models not only follow instruction reliably but also maintain an understanding on task objectives, unlike OM models that narrowly focus on the problem-solving aspect.
5.2 Keypoint-based Explanation Generation
Metric
We assess explanation quality using three criteria: (1) Correctness – whether the model’s reasoning is logically sound and leads to a valid solution; (2) Fidelity – whether the explanation aligns with the reasoning and intent of the reference, regardless of surface form; (3) Visual Referencing – whether the explanation refers to the same key visual components (e.g., points, lines, angles, shapes) as the reference. Each criterion is rated on a 5-point Likert scale. We report annotation results from both human evaluators (Zheng et al., 2023) and an LLM-based evaluator (GPT‑4o (Achiam et al., 2023)). Complementary results based on conventional text similarity metrics (Papineni et al., 2002; Banerjee & Lavie, 2005) are provided in Appendix A.
GPT-Eval Human-Eval
Model Correctness Fidelity Referencing Correctness Fidelity Referencing
OG Molmo 2.366 1.907 2.548 - - -
LLaVA-1.6 1.822 1.549 1.973 - - -
[…]
OM Math-PUMA 1.775 1.614 1.588 1.036 1.036 1.030
URSA 1.150 1.142 1.157 - - -
CG Gemini 2.0 Flash 3.872 3.508 4.126 4.361 4.060 4.084
[…]
Table 4: Experimental results for the Keypoint-based Explanation Generation task on MathExplain. Correctness, fidelity, and referencing are rated on a 1–5 Likert scale by both GPT and human evaluators.
Result
Table 4 shows that stronger, general-purpose models consistently outperform smaller, math-specific baselines, continuing the performance trend observed in the previous task. According to GPT-Eval, the closed-source model Gemini 2.0 Flash clearly leads across all three dimensions, Correctness, Fidelity, and Referencing. Among the open-source models, Qwen2.5-VL performs notably well, ranking third overall in GPT-Eval. In contrast, the math-specialist models struggle to produce logically consistent explanations, as reflected in their lower scores across all three dimensions. This underperformance is particularly striking for URSA, which previously excelled in mathematical benchmarks (Luo et al., 2025). Our qualitative analysis confirms that the specialist are optimized for problem solving rather than solution explaining. Human-Eval reveals similar findings. Gemini 2.0 Flash again receives the highest ratings, while Qwen2.5-VL continues to moderately outperform Math-PUMA. Overall, both GPT-Eval and Human-Eval reveal consistent trends—Gemini 2.0 Flash demonstrates a strong capacity to generate effective keypoint-based explanations, while most open-source models, especially those in the OM category, lag significantly behind. This highlights the performance gap between top-performing closed-source models and their open-source counterparts in visual solution-explaining tasks.
GPT-Eval
Model Correctness Fidelity Referencing
[…]
Math-PUMA 1.233 1.233 1.216
Gemini 2.0 Flash 3.749 3.343 3.992
Geometry Graph
MCQA SA Overall MCQA SA Overall
[…]
0.682 0.488 0.607 0.695 0.649 0.677
Table 5: Results of the Keypoint-based Explanation Generation task under the ablated (no-answer) setting (left), along with problem-solving performance under the same condition (right).
5.3 Does explanation quality reflect problem-solving ability?
Given that recent open-source MLLMs are generally believed to possess reasonable problem-solving capabilities, our findings suggest that explanation generation remains a distinct and more challenging capacity. To investigate this hypothesis, we examine the relationship between explanation quality and problem-solving performance.
For this analysis, ground-truth answers are withheld during inference. Models are first prompted to generate an explanation without access to the correct answer (“blind explanation”), and subsequently asked to produce a hypothetical answer based solely on their self-generated explanation.
Result
Table 5 (left) presents GPT-based evaluations (GPT-Eval) under the no-answer condition for the Keypoint-based Explanation Generation task, in which models are required to produce explanatory content without access to the final answer. Despite this constraint, performance patterns remain broadly consistent with those observed in previous experiments. However, the absolute scores drop across all models, indicating that the complexity of explanation generation hinges on the availability of the correct answer.
Table 5 (right) shows the accuracy of the answers that the models generated conditioned on their own explanations. Gemini 2.0 Flash, which previously demonstrated strong performance in visual solution-explaining, consistently outperforms other models across both Geometry and Graph tasks, achieving notably high accuracy in both MCQA and SA settings. In contrast, Math-PUMA performs poorly not only in Keypoint-based Explanation Generation but also in problem-solving accuracy. Notably, it is even outperformed by Qwen—a model not specifically optimized for mathematics—suggesting that the ability to generate coherent explanations may be linked to stronger problem-solving performance.
6 Conclusion
We introduce visual solution-explaining, a task that requires essential mathematics-teaching capabilities for delivering intuitive and insightful explanations alongside problem-solving. We also propose MathExplain, which assesses visual solution-explaining performance through two subtasks: (1) Visual Keypoint Identification, challenging models to generate visual keypoints that are crucial for subsequent explanations, and (2) Keypoint-based Explanation Generation, requiring models to integrate identified keypoints into coherent, solution-driven narratives. Our experimental results show that open-source models struggle to accurately identify the key visual elements necessary for effective problem comprehension. This shortcoming is consistently reflected in their explanatory text generation performance, particularly in mathematics-oriented models fine-tuned on domain-specific data. We believe that visual solution-explaining and MathExplain will inspire further research on developing models for educational applications, ultimately paving the way for real-world mentors capable of providing clear, student-friendly solutions. All code and datasets will be made publicly available.
7 Limitations
While visual solution-explaining has demonstrated considerable promise in handling Visual Keypoint Identification through a combination of parser and matching modules, the Keypoint-based Explanation Generation component introduces additional open-ended complexities. Specifically, it is more challenging to provide direct, high-level evaluations in situations that demand nuanced, structured outputs. Current matching techniques are not yet fully equipped to capture the broader semantic scope required for such tasks. Nevertheless, these challenges present exciting opportunities for ongoing refinement. As models evolve to produce more robust structured outputs, we anticipate that the evaluation and matching procedures will similarly advance, paving the way for even more comprehensive and intuitive assessment methods.
8 Ethical Considerations
In this paper, we introduce the MathExplain benchmark, which is curated from an in-house mathematics education platform [^1] https://mathpresso.com/en. During the data collection process, all materials were carefully curated to ensure full compliance with copyright regulations. For data annotation, we employed individuals who are proficient in English and hold undergraduate degrees, ensuring that they possess adequate mathematical problem-solving skills.
Appendix A Experiments Details
In this study, experiments were conducted using the LMMs-eval repository (Li et al., 2024). This repository provides a comprehensive framework for evaluating multi-modal models across various tasks. In all GPT evaluations, we used the gpt-4o-2024-08-06 endpoint. When conducting experiments with open-source multimodal models, we leveraged the official implementation codes in conjunction with publicly available weights from the Huggingface Hub  [^2] https://huggingface.co/models. The following model parameters were used for each model:
  • Math-PUMA: Math-PUMA/Math-PUMA_Qwen2VL-7B
  • URSA: URSA-MATH/URSA-RM-8B
  • llava-1.6: llava-hf/llava-v1.6-mistral-7b-hf
  • Qwen2.5-VL: Qwen/Qwen2.5-VL-7B-Instruct
  • Molmo: allenai/Molmo-7B-D-0924
These models were evaluated in our benchmark, which included tasks designed to assess both visual understanding and textual explanation capabilities. The selection of models spans a range of architectures and performance levels, providing insights into current advancements in multi-modal learning.
A.1 Computational Resources
For closed-source models, we used the OpenAI API and Gemini Developer API to infer the output of GPT-4o and Gemini 2.0. For open-source models such as Math-PUMA, URSA, LLaVA, QWEN-2.5-VL, and Molmo, inference was performed using a NVIDIA A6000 48GB GPU. While the exact inference speed varies depending on the task, model, and lengths of prompts and responses, a query takes about 50 seconds to be answered.
A.2 Evaluation on automatic metrics
Model #Params BLEU-2 BLEU-4 ROUGE-L METEOR BertScore
[…]
LLaVA-1.6 7B 0.130 0.059 0.175 0.286 0.835
[…]
OM Math-PUMA 7B 0.006 0.002 0.118 0.058 0.818
URSA 8B 0.020 0.006 0.078 0.075 0.735
[…]
GPT-4o - 0.095 0.045 0.160 0.300 0.814
Table 6: Automated Evaluation Metrics for the Generating Explanation Text Task on MathExplain
Appendix B Benchmark Details
Our MathExplain consists of a total of 17 chapters and 33 sections as shown in Table 7
Chapter Title Section Title
Basics of Geometry Basic Geometric Figures
Construction and Congruence
Coordinate Plane and Graphs Coordinate Plane and Graphs
Differential Calculus Differentiation of Various Functions
Differentiation Derivative and Derivative Function
Applications of Derivatives
Equations and Inequalities Quadratic Equations and Functions
Equations of Geometric Figures Transformations of Figures
Equation of a Circle
Coordinate Plane
Equations of Straight Lines
Exponential and Logarithmic Functions Exponential and Logarithmic Functions
Functions Linear Functions and Their Graphs
Functions
Rational and Irrational Functions
Relationship Between Linear Functions and Equations
Integral Calculus Applications of Definite Integrals
Various Integration Techniques
Integration Indefinite and Definite Integrals
Applications of Definite Integrals
Plane and Solid Figures Properties of Solid Figures
Properties of Plane Figures
Properties of Circles Circle and Line
Inscribed Angles
Properties of Figures Properties of Quadrilaterals
Properties of Triangles
Quadratic Functions Graph of the Quadratic Function $y=ax^{2}+bx+c$
Similarity and the Pythagorean Theorem Pythagorean Theorem
Similarity of Figures
Applications of Similarity
Trigonometric Functions Meaning and Graphs of Trigonometric Functions
Law of Sines and Law of Cosines
Trigonometric Ratios Trigonometric Ratios
Applications of Trigonometric Ratios
Table 7: Overview of Chapters and Sections in the Dataset. This table presents the 17 chapters and their corresponding 33 sections covered in our dataset.
Appendix C Prompts
This section compiles all the prompts used in our experiments. The prompts in Figure 6 and Figure 7 are used to elicit model outputs for the visual keypoint generation and explanatory text generation tasks, respectively. For the visual keypoint generation task, since model outputs are not always in JSON format, we applied a GPT-based parsing step with the prompt shown in Figure 8. Model responses are evaluated using the prompts in Figure 9 for explanatory text generation task and Figure 10 for visual keypoint generation task.
Q: As shown in the figure, there are 5 points: A, B, C, D, and E. When selecting two points among them to form straight lines and rays, let the number of straight lines be a and rays be b. Find the value of a + b. ### Answer: 19 ### You are a teacher, who helps students to solve math problems. You have to solve the following problem and explain the solution to your students. For image captioning, you should explain what should be changed in the given image to help students understand your explanation better. Additionally, this problem is a problem of Functions/Linear Functions and Their Graphs chapter. You should explain the problem in the context of the chapter and section. Your output must strictly follow the JSON format below with no additional text or commentary. ### OUTPUT Example: { {”element”: ”Point C”, ”description”: ”A point located at the right end of the line”},
   {”element”: ”Line AB”, ”description”: ”A line connecting point A to point B”}
Figure 6: The prompt used for eliciting model-generated visual keypoints in the MathExplain benchmark. It is designed to obtain response in the form of visual keypoints to a given math problem. The inputs of the prompts are boldfaced.
Q: As shown in the figure, there are 5 points: A, B, C, D, and E. When selecting two points among them to form straight lines and rays, let the number of straight lines be a and rays be b. Find the value of a + b. ### Answer: 19 ### ### Difference between the original image and the solution image ### Line AE: A straight line connecting points A and E is shown
Line BE: A straight line connecting points B and E is shown
Line CE: A straight line connecting points C and E is shown
Line DE: A straight line connecting points D and E is shown
You are a math teacher helping students understand how to solve problems clearly and effectively. Given a problem description, problem image and a list of key elements introduced or highlighted in the solution image, write an educational explanation that helps students. Additionally, this problem is a problem of Functions/Linear Functions and Their Graphs chapter. You should explain the problem in the context of the chapter and section. Make sure to reference both the original components from the problem image and any new annotations, highlights, or added elements from the solution image to enhance understanding. ### OUTPUT Example: {
solution_text:
Figure 7: The prompt used for eliciting explanatory text generation task in MathExplain benchmark. It is designed to generate educationally effective explanations for the given math problem. The inputs of the prompts are boldfaced.
You are a helpful assistant that can parse the text into a JSON format.
- JSON: If the input is in a JSON-like format but contains formatting errors, extract the relevant pieces of information and convert them into a clean, valid JSON structure as described below.
- Plain Text: If the input is in a descriptive, non-JSON format, extract the necessary elements and format them as a JSON array using the structure defined below.
## JSON Output Requirements:
- The output must be a valid JSON array.
- Each item in the array must be an object containing exactly two keys: ”element” and ”description”.
- No two items should refer to the same element. Each element must be uniquely represented.
- The ”element” key must use only one of the allowed formats listed below:
   - Points:
      - Point A, Point (0,7)
   - Lines:
      - Line AB, Straight line AB, Pointed line AB
   - Angles:
      - Angle ABC, Angle 30°, Angle A
   - Regions:
      - Region S_1
   - Symbols:
      - Right angle mark N
      - Parallel mark N
      - Equal length mark N
      - Equal angle mark N (use this when two or more newly shown angles are equal. If only one angle is shown, use Angle instead.)
      - Length label N
### Each object must contain:
- ”element”: the name or title of the element (must match one of the accepted formats above)
- ”description”: a brief explanation of what the element represents or describes
### Example Output Format:
“‘json
   {”element”: ”Point (0,7)”,
   ”description”: ”Shown as the y-coordinate of point B”},
   {”element”: ”Line BE”,
   ”description”: ”A line connecting points B and E is shown”},
   {”element”: ”Region S_1”,
   ”description”: ”A shaded area representing the positive values of velocity v(t) over the interval 0 ≤ t ≤ 2”},
   {”element”: ”Angle x”,
   ”description”: ”Represents that angles EAG, EGA, FDG, and FGD are equal”},
   {”element”: ”Angle 34°”,
   ”description”: ”Indicates the size of angle EGF”},
   {”element”: ”Equal length mark 1”,
   ”description”: ”Indicates that line segments KH and KG are equal in length”},
   {”element”: ”Equal length mark 2”,
   ”description”: ”Indicates that line segments EJ and JF are equal in length”},
   {”element”: ”Equal angle mark 1”,
   ”description”: ”Indicates that angles JEF, JFE, and EFB are equal”},
   {”element”: ”Equal angle mark 2”,
   ”description”: ”Indicates that angles KGH, KG, and HGC are equal”},
   {”element”: ”Parallel mark”,
   ”description”: ”Indicates that lines DE and BC are parallel”},
   {”element”: ”Length label 1”,
   ”description”: ”Line DE is labeled with a length of 10”},
   {”element”: ”Length label 2”,
   ”description”: ”The distance between points B and C is labeled as 20”},
   {”element”: ”Right angle mark 1”,
   ”description”: ”Indicates that angle OAP is a right angle”},
   {”element”: ”Right angle mark 2”,
   ”description”: ”Indicates that angle APO is a right angle”}
### Your Task:
Identify the key elements and construct the correct JSON array with ”element” and ”description” fields.
IMPORTANT: Your output must not include any informations not about the visual keypoints.
For example, even if the input contains ”Chain rule”, you should not include it in the output because it is not a visual keypoint.
[…]
## Input:
{”problem_description”: ”In the given image, there are five points: A, B, C, D, and E”,}
{”image_description”: ”The image shows a diagram with five points: A, B, C, D, and E. Points A and B are connected by a straight line, and points B and C are connected by another straight line. Points C and D are connected by a straight line, and points D and E are connected by another straight line. There are also two rays: one from point A to point C and another from point D to point E. The diagram is a simple representation of the problem.”,
{”solution”: ”To find the value of a + b, we need to count the number of straight lines and rays in the diagram. There are three straight lines: AB, CD, and DE. There are also two rays: AC and AD. Therefore, the value of a + b is 3 (straight lines) + 2 (rays) = 5.”
Figure 8: Prompt for parsing model outputs to ensure error-free, valid JSON formatting during the visual keypoint generation task in MathExplain. The inputs of the prompts are boldfaced.
You are evaluating the quality of an AI-generated explanation for a math problem involving geometry or graph-based reasoning.
You will be given two texts:
1. A reference explanation written by a human teacher.
2. An AI-generated explanation written by a model.
Your task is to compare the two explanations and assess how accurately and effectively the AI-generated explanation captures the key geometric concepts and reasoning presented in the reference.
Please evaluate the model’s explanation and provide four scores based on the criteria below:
### Scoring Criteria
1. Correctness
- Does the reasoning presented by the model make sense and help solve the problem appropriately?
- Rate on a Likert scale: **1, 2, 3, 4, or 5**
2. Reference Alignment
- Does the model follow the same logical reasoning and intent as the reference explanation, even if the wording differs?
- Rate on a Likert scale: **1, 2, 3, 4, or 5**
3. Use of Key Visual Elements
- Does the AI explanation refer to the same critical visual components (e.g., points, lines, angles, shapes) as the reference?
- Alternative terminology is acceptable if it clearly refers to the same element or serves the same purpose.
- Rate on a Likert scale: **1, 2, 3, 4, or 5**
4. Mathematical Accuracy
- Does the model produce the same final answer as the reference?
- If the model’s answer is correct, give a score of **1**.
- If the model’s answer is incorrect, give a score of **0**.
### Output Format
Important: Report your rating using the exact format below:
Rating: [[x, y, z, w]]
— where ‘x‘ is your score for correctness, ‘y‘ for reference alignment, ‘z‘ for use of visual elements, and ‘w‘ for mathematical accuracy.
Figure 9: GPT Prompt for evaluating model output on explanatory text generation task in MathExplain. The inputs of the prompts are boldfaced.
You are a careful and precise visual reasoning assistant. Your job is to analyze and compare diagram annotations.
You will be given:
1. A Reference JSON that describes the correct elements and transformations in a diagram (from a problem image to a solution image).
2. An AI-Generated JSON that contains elements proposed by an AI model.
Your task is to compare the Reference JSON with the AI-Generated JSON to evaluate how accurately the AI has captured the intended diagram changes.
You should understand the transformation from the problem image to the solution image and assess whether the AI-generated elements correctly reflect those changes (Reference elements).
Element labels do not need to match exactly; if their meaning, location, or role in the diagram is the same or very similar, they should be considered as a match.
## Matching Criteria:
- If an AI-generated element performs the same role or function in the diagram as a reference element (even with a different label), you should map them together.
- One reference element can match one AI element. If no match is found, leave the generated side null.
- If the AI-generated JSON contains duplicate or redundant entries (e.g., different keys with identical descriptions), select the first one.
## Output Format
For each reference element, output a single object in the following format:
“‘json
      { ”Reference”: ”Reference Element”, ”Generated”: ”Matched AI Element” }
If there is no meaningful match for a reference element, set ‘”Generated”‘ to null.
IMPORTANT: Make sure that your output contain same number of elements as the reference JSON.
IMPORTANT: All results should be returned as a JSON array.
## Example
### Reference JSON
      {”element”: ”Point C”, ”description”: ”A point located at the right end of the line” },
      {”element”: ”Line AB”, ”description”: ”A line connecting point A to point B” },
      {”element”: ”Angle ABC”, ”description”: ”The angle formed by connecting the three points” },
      {”element”: ”Right angle symbol”, ”description”: ”A right angle marker at the intersection of two lines X and C” },
      {”element”: ”Parallel mark”, ”description”: ”Indicates that lines DE and BC are parallel” }
### AI-Generated JSON
      {”element”: ”Point d”, ”description”: ”A point located at the right end of the line” },
      {”element”: ”Angle ABD”, ”description”: ”An angle marked using the newly added segment AB” },
      {”element”: ”x angle symbol”, ”description”: ”An angle marked below by connecting the center of line AB to point D” },
      {”element”: ”x angle”, ”description”: ”An angle marked below by connecting the center of line AB to point D” },
      {”element”: ”c angle”, ”description”: ”Indicates that angle XCA is a right angle” }
### Output
“‘json[
      {”Reference”: ”Point C”, ”Generated”: ”Point d” },
      {”Reference”: ”Line AB”, ”Generated”: ”Angle ABD” },
      {”Reference”: ”Angle ABC”, ”Generated”: ”Angle ABD” },
      {”Reference”: ”Right angle symbol”, ”Generated”: ”c angle” },
      {”Reference”: ”Parallel mark”, ”Generated”: ”null” }
### Reference JSON
{”element”:”Line AE”, ”Generated”: ”A straight line connecting points A and E is shown”}
      {”element”:”Line BE”, ”Generated”: ”A straight line connecting points B and E is shown”}
      {”element”:”Line CE”, ”Generated”: ”A straight line connecting points C and E is shown”}
      {”element”:”Line DE”, ”Generated”: ”A straight line connecting points D and E is shown”}
### AI-Generated JSON
{”element”: ”Point A”, ”description”: ”One of the five points shown in the diagram”}
      {”element”: ”Point B”, ”description”: ”One of the five points shown in the diagram”}
      {”element”: ”Point C”, ”description”: ”One of the five points shown in the diagram”}
      {”element”: ”Point D”, ”description”: ”One of the five points shown in the diagram”}
      {”element”: ”Point E”, ”description”: ”One of the five points shown in the diagram”}
      {”element”: ”Line AB”, ”description”: ”A straight line connecting points A and B”}
      {”element”: ”Line BC”, ”description”: ”A straight line connecting points B and C”}
      {”element”: ”Line CD”, ”description”: ”A straight line connecting points C and D”}
      {”element”: ”Line DE”, ”description”: ”A straight line connecting points D and E”}
      {”element”: ”Ray AC”, ”description”: ”A ray starting from point A and passing through point C”}
      {”element”: ”Ray AD”, ”description”: ”A ray starting from point A and passing through point D”}
Figure 10: GPT Prompt for evaluating model output on visual keypoint generation task in MathExplain. The inputs of the prompts are boldfaced.